Grade 12Mechanics

A body weighs 63N on the surface of the earth. The gravitational force on it due to the earth at a height equal to half of the radius of the
earth is 7xN, then x=---

suyash sunil rokade
8 Years agoGrade 12

4 Answers

Arun
ApprovedApproved Tutor Answer8 Years ago
Weight of the body, W = 63 N
Acceleration due to gravity at height h from the Earth’s surface is given by the relation:
g‘ = g / [1 + ( h / Re) ]2
Where,
g = Acceleration due to gravity on the Earth’s surface
Re = Radius of the Earth
For h = Re / 2
g‘ = g / [(1 + (Re / 2Re) ]2
g / [1 + (1/2) ]2  =  (4/9)g
Weight of a body of mass m at height h is given as:
W‘ = mg
m × (4/9)g  =  (4/9)mg
= (4/9)W
= (4/9) × 63  =  28 N.
Hafiz shah s.r
ApprovedApproved Tutor Answer8 Years ago
Weight on the surface of earth=63N
Mg=63N
On the surface of earth gravitational force = Mg=63N
M=63/g.
 
Gravitational force at a ht  ,H =R/2=
Mg¹.
g¹=g{R²/(R+H)²}...on solving g¹=4g/9
:-Mg¹=M×4g/9
Where M=63/g.
=63/g×4g/9=7×4=28N
:-. X= 4.
Hope you understood....
 
Nabin
8 Years ago
Weight on earth surface=63N
Mass of the body=6.3(g=10)
height of the body=radius/2
now,
g1={r^2/(r+h)^2}g
     =40/9
Then
Force=m*g1
7x=6.3*40/9
x=4
 
Abhinav Vaisoya
7 Years ago
We know that w=mg
therefore w=f
f=48n
Or
48=\frac{GMm}{r^{2}}
When distance from surface of the Earth is 6400,
r becomes doubled so,
new weight=\frac{GMm}{2r^{2}}
mg2=\frac{GMm}{4r^{2}}
mg2=\frac{GMm}{r^{2}}*\frac{1}{4}
mg2=48*\frac{1}{4}
mg2=12
So new weight will be 12N 
Hope this helps you!